# Static Spherical Vacuum Solution to Bumblebee Gravity with Time-like VEVs

### v3

@@ V.2 Violation of the weak cosmic censorship conjecture
The weak cosmic censorship conjecture (WCCC), proposed by Sir Roger Penrose [16], serves as a method to save the predictability of general relativity [17]. However, our solutions in this work provide explicit examples that counter this conjecture. In fact, the solution Eq. (23) and the corresponding Kretschmann scalar Eq. (24) indicate that there is a singularity located at $r=0$, but no horizon exists in this case, meaning the existence of a naked singularity, and the violation of the WCCC consequently. Nonetheless, we have argued that the solutions in Sec. IV are unstable, therefore the violation of the WCCC appears to be harmless, since the solutions would finally decay into the flat space-time.
V.3 Summary
In this work, we study the static spherical vacuum solution for bumblebee gravity with time-like VEVs as $b_{\mu}=(b_{t}(r),0,0,0)$. The results suggest that for general VEVs, there is no consistent non-trivial solution if $b\neq\sqrt{2/\kappa}$. With $b=\sqrt{2/\kappa}$, we find two non-trivial solutions as Eq. (23) and Eq. (41), none of which is Schwarzschild-like. And we argue that these solutions are unstable, which comes to the rescue when the predictability of the theory is ruined because of the existence of naked singularities. Since there are no Schwarzschild-like solutions for bumblebee gravity with time-like VEVs, and only the flat space-time solution is stable, we probably do not need to consider the case of time-like VEVs when related to real physics in future researches. However, the newly obtained solutions give rise to further questions, such as what is the reason for the existence of a solution when $b=\sqrt{2/\kappa}\simeq E_{Pl}$.
Acknowledgements.
@@
The authors are grateful for valuable discussions with Hanlin Song and Peixiang Ji. This work was supported in part by the National Science Foundation of China under Grant No. 12347101. HL was also supported by the start-up fund of Chongqing University under No. 0233005203009, and JZ was supported by the start-up fund of Chongqing University under No. 0233005203006.

### v5

@@ 5.2 Violation of the weak cosmic censorship conjecture
The weak cosmic censorship conjecture (WCCC), proposed by Roger Penrose [32], serves as a method to save the predictability of general relativity [33]. Obviously, our solutions in this work provide explicit counterexamples to this conjecture, as we will show later. Nonetheless, we have argued that the solutions in Sec. 4 are unstable; therefore, the violation of the WCCC appears to be harmless, since the solutions would finally decay into the flat space-time. And the natural expectation of the value of $\ell$ also supports this argument.
Regardless of the stability of the solutions, we analyze the properties of the naked singularities in more detail. The question is whether the naked singularities in our solutions are global, where a singularity is globally naked if there is a future-directed causal curve with one end on the singularity and the other end on the future null infinity [25, 34]. Although the best way to secure the prediction power is cover the singularity behind an event horizon, a singularity could still be allowed if it is not globally naked or even not locally naked. Let us first consider the solution Eq. (28) and only focus on the radial part of this metric. Then, for a outward-directed photon, its geodesic motion is governed by
$0=-\left(\frac{r}{R_{0}}\right)^{\frac{2(2-\ell)}{\ell}}dt^{2}+\frac{4}{\ell^{2}}dr^{2},$ (56)
where the parameter $\ell$ is restricted by $0\leq\ell\leq 2$. A straightforward calculation gives the propagating time taken by a photon which is emitted at $r=0$ moving to $r=R>0$ as
$\Delta t=\lim_{\epsilon\to 0^{+}}\int_{\epsilon}^{R}\frac{2}{\ell}\left(\frac{r}{R_{0}}\right)^{\frac{\ell-2}{\ell}}=\lim_{\epsilon\to 0^{+}}\left.\frac{r(r/R_{0})^{\frac{l-2}{\ell}}}{\ell-1}\right|^{R}_{\epsilon}.$ (57)
If $\ell>1$ then $\Delta t$ is finite, and the singularity is globally naked; while for $\ell<1$ we have $\Delta t=\infty$, making the singularity even not naked locally. Next, we consider the radial motion of a photon in the spacetime determined by Eq. (46). Similarly, considering a photon emitted outwards at $r=R_{s}$ and reach an observer at $r=R>R_{s}$, we can write the propagating time as
$\Delta t=\lim_{\epsilon\to 0^{+}}\int_{R_{s}+\epsilon}^{R}\left(1-\frac{R_{s}}{\rho}\right)^{\ell-2}d\rho,$ (58)
where we again require $0\leq\ell\leq 2$ such that energy conditions are satisfied. By estimating this integral, we come to the same conclusion that for $\ell>1$ there is a globally naked singularity, while for $\ell<1$ the singularity is not naked, even locally.
A few comments are in order. The parameter $\ell=\xi b^{2}$ plays a very important role in the analysis of the singularities. As we can see, regardless of which solutions we are considering, $\ell>1$ always means globally visible singularities, giving counterexamples of the weak cosmic censorship conjecture. On the other hand, however, if $\ell<1$, even though the singularities are not hidden by event horizons, our ability to make predictions is still maintained. Meanwhile, no energy condition is violated for both cases. But which case is preferred? Let us come back to the action Eq. (1), which contains not only the usual gravity and matter contributions, but also a vector field sector non-minimally coupled to gravity. The coupling constant of the non-minimally coupled term can be read off as $\xi/2\kappa=\ell/4$, where we have used $\ell=\xi b^{2}=2\xi/\kappa$. Since $\ell$ is effectively the degree to which the vector field is coupled to gravity non-minimally, and generally we expect that non-minimal couplings are small, $0\sim\ell<1$ is naturally favored  [^6] Besides, in Ref. [9], the authors showed that this parameter has an upper-bound of $\ell<10^{-13}$. Although the solution in Ref. [9] was obtained with a space-like background VEV, the parameter $\ell$ in both cases should have the same order of magnitude.. Consequently, the singularities in our solutions bring no problem to the power of making predictions.
It is also beneficial to examine the geodesic completeness next. Specifically, let us consider radial time-like geodesics
$d\tau^{2}=g_{tt}dt^{2}+g_{rr}dr^{2},$ (59)
where $\tau$ is the proper time along the geodesics. Since the solutions to be examined always have a time-like Killing vector $K^{\mu}=(\partial_{t})^{\mu}=\delta^{\mu}_{\ 0}$, we can construct a constant as
$E=-K_{\mu}\frac{dx^{\mu}}{d\tau}=-g_{tt}\frac{dt}{d\tau}.$ (60)
Combining Eq. (59) and Eq. (60), we obtain
$\frac{d\tau}{dr}=-\frac{\sqrt{-g_{tt}g_{rr}}}{\sqrt{E^{2}+g_{tt}}}.$ (61)
First we check the geodesic completeness for metric (28), for which Eq. (61) leads to
$\tau=-\frac{2}{\ell}\int_{R}^{0}\frac{dr}{\sqrt{E^{2}(r/R_{0})^{2(\ell-2)/\ell}-1}},$ (62)
where the arbitrary and finite radius $R>0$. We do not need to explicitly perform the integration, and a qualitative analysis shows that this solution is geodesic incomplete. Actually, there is a critical radius for each choice of $E$: $r_{c}=E^{\ell/(2-\ell)}R_{0}$. If $R<r_{c}$, given $0<\ell<2$, it is straightforward to show that the integration (62) is indeed finite, indicating geodesic incompleteness. From the analysis of time-like geodesics, a rather weird but intriguing behavior of this spacetime emerges: to go to infinity $r=+\infty$, a massive test particle should have infinitely large energy, which can be seen from $r_{c}=E^{\ell/(2-\ell)}R_{0}$. In this sense, the spacetime is even closed for massive particles, and interestingly, this spacetime appears to exhibit a nature of confinement. Similarly, for the metric (46), we have
$\tau$ $=-\int_{R}^{R_{s}}\frac{dr}{\sqrt{E^{2}-(1-R_{s}/r)^{2-\ell}}}$
$=-R_{s}\int_{R/R_{s}}^{1}\frac{dr}{\sqrt{E^{2}-(1-1/r)^{2-\ell}}},$ (63)
noting that the singularity is located at $r=R_{s}$ rather than $r=0$. Since again we require $0<\ell<2$, the proper time could be finite with suitable choices of the ratio $R/R_{s}$ and the “energy” $E$. As a result, this solution is also geodesic incomplete. But in this spacetime, a massive test particle can arrive at infinity as long as $E>1$, very similar to the case in the Schwarzschild spacetime.
5.3 Photon spheres and observations
Although the solutions do not have clear physical pictures, i.e., whether they are merely mathematically acceptable solutions, or are results of some physical processes such as gravitational collapse, it is helpful to study their aspects of observations.
Here we would like to analyze the photon spheres of solutions Eq. (28) and Eq. (46). A photon sphere is defined to be a nowhere-spacelike hypersurface, such that any photon whose 4-velocity is initially tangent to the surface will always remain in the surface if not altered [35]. More intuitively, as a light ray approaches the photon sphere, the Einstein bending angle tends to be infinitely large [36, 37]. For a Schwarzschild black hole, the gravitational lensing would cause relativistic images due to photon spheres, besides the ordinary primary and secondary images [36]. Furthermore, the existence or non-existence of photon spheres around a singularity (naked or not) has important implications for gravitational lensing, which can be compared to realistic observations. In the literature, a singularity with or without photon spheres is termed a weakly or strongly naked singularity (WNS or SNS) respectively. To determine the photon spheres, considering a general static spherically symmetric spacetime described by [37, 38]
$ds^{2}=-B(r)dt^{2}+A(r)dr^{2}+D(r)r^{2}d\Omega^{2},$ (64)
we can obtain the photon sphere by solving
$2D(r)B(r)+rD^{\prime}(r)B(r)-rB^{\prime}(r)D(r)=0,$ (65)
where only constant $r$ solutions are considered. As an example, the photon sphere obtained by solving Eq. (65) for a Schwarzschild metric is $r=3M$, which is exactly the desired result.
First, we examine the spacetime expressed by Eq. (28). Combining Eq. (28) with Eq. (65), we obtain
$\frac{4(1-\ell)r^{4/\ell-2}}{\ell R_{0}^{2-4/\ell}}=0,$ (66)
which has no reasonable solution since $0<\ell<2$. Thus, we conclude that the singularity of Eq. (28) is strongly naked. Then we perform the same analysis for the spacetime Eq. (46), and Eq. (65) gives
$\frac{2(r-R_{s})\left(r-(2-\ell)R_{s}\right)}{r^{2}}=0,$ (67)
leading to $r=R_{s}$ or $r=(2-\ell)R_{s}$. But we have already shown that $R_{s}$ is actually a singularity, so there is only one photon sphere located at $r=(2-\ell)R_{s}$. Furthermore, we also argue that, to make the singularity globally (even locally) invisible, $\ell$ should be less than $1$, making $(2-\ell)R_{s}>R_{s}$, and rendering the singularity weakly naked. Put differently, requiring the singularity to be weakly naked in this case is equivalent to having a photon sphere surrounding the singularity. According to Refs. [37, 38], a Schwarzschild black hole and a WNS share the same qualitative gravitational lensing features, and the quantitative values are also very close; while for a SNS quite different qualitative features emerge, distinguishable from a Schwarzschild black hole. As a result, the type of spacetime expressed by Eq. (46) cannot be distinguished from a Schwarzschild black hole by gravitational lensing, but the gravitational lensing can be used to distinguish Eq. (28) from a Schwarzschild black hole.
6 Summary
In this work, we study the Class I ($b_{r}\equiv 0$) static spherical vacuum solution for bumblebee gravity with time-like VEVs as $b_{\mu}=(b_{t}(r),0,0,0)$ such that $b^{\mu}b_{\mu}=-b^{2}=\text{const}$. Different from the Class II ($R_{rr}\equiv 0$) solutions in Ref. [23], the results suggest that for general VEVs, there is no consistent non-trivial solution if $b\neq\sqrt{2/\kappa}$. With $b=\sqrt{2/\kappa}$, we find two non-trivial solutions as Eq. (28) and Eq. (46), none of which is Schwarzschild-like, but the second one shares some similarity with Janis-Newman-Winicour spacetime [24]. The solutions both exhibit naked singularities, but we argue that these solutions are unstable, which seems to make the singularities harmless. Nevertheless, we show that in the minimal coupling limit, i.e., $\ell\propto\xi b^{2}\ll 1$, the singularities are not globally visible, and they are even locally invisible, thus the weak cosmic censorship conjecture appears to be still valid. Then we analyze the photon spheres and find that Eq. (28) has no photon spheres, while Eq. (46) has a photon sphere outside the singularities. As a result, the gravitational lensing features of Eq. (28) could be very different from those of a Schwarzschild black hole; however, by only utilizing gravitational lensing features, Eq. (46) cannot be distinguished from a Schwarzschild black hole. Finally, the newly obtained solutions give rise to further questions, such as what is the reason for the existence of a solution when $b=\sqrt{2/\kappa}\simeq E_{Pl}$, and we hope future studies could shed light on the questions.
As a final remark, this study has primarily examined the theoretical properties of the solutions arising in the bumblebee model with time-like VEVs. A distinctive feature of our results is that the geometries obtained correspond to naked singularity solutions rather than conventional black holes. This distinction suggests that their dynamical and observational signatures—such as quasinormal modes and greybody factors—may differ substantially from those typically associated with black-hole spacetimes [39]. A systematic computation of these quantities [19, 39] would therefore deepen our understanding of the physical behavior of these solutions and, more broadly, enrich the study of naked singularities by providing a new and explicitly solvable theoretical example. In addition to dynamical probes, it would also be worthwhile to explore the implications of these solutions for gravitational lensing and the formation of shadow-like structures. Since the causal and geometric properties of naked singularity spacetimes differ markedly from those of black holes, their lensing patterns and shadow profiles may exhibit qualitatively new features, opening up novel possibilities for observational discrimination [37, 38]. Such analyses could facilitate more direct comparison with astrophysical data, including the Event Horizon Telescope observations of M87* and Sgr A*. While these extensions would significantly strengthen the observational relevance of our results, they lie beyond the scope of the present work and are left for future investigation.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
@@
The authors are grateful for the valuable discussions with Hanlin Song and Peixiang Ji. We also thank Kumar Shwetketu Virbhadra for his valuable correspondence. This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12547101. HL was also supported by the start-up fund of Chongqing University under No. 0233005203009, and JZ was supported by the start-up fund of Chongqing University under No. 0233005203006.